3 years older than four times the age of a younger person. The sum of their ages is 13. Find their ages


Answer by ikleyn(52803)

.

x + (4x+3) = 13

    5x     = 13-3

    5x     = 10

     x     = 10/5 = 2.


ANSWER.  2 years and 2*4+3 = 11 years.


CHECK.  2 + 11 = 13 years total.   ! Correct !

Solved.